Will someone please help me solve these problems.
1) Find the greatest common factor of the two expression: Simplify
20uy^8z^6 and 28u^7y^4
2) Find the least common multiple of the two expressions: Simplify
3t^6x^8y^5 and 12x^3y^6
3) Factor Completely: 27z^3+125
4) Factor: 9t+wt+9y+yw
5) Factor: -8t+xt-24x+3x^2


Found 2 solutions by stanbon, nyc_function:
Answer by stanbon(75887)

1) Find the greatest common factor of the two expression: Simplify
20uy^8z^6 and 28u^7y^4
---
gcf for 20 and 28 is 4
gcf for u and u^7 is u
gcf for y^8 and y^4 is y^4
gcf for z^6 and no z is no z
---------------------------------
GCF = 4uy^4
========================================================================
2) Find the least common multiple of the two expressions: Simplify
3t^6x^8y^5 and 12x^3y^6
lcm of 3 and 12 is 12
lcm of t^6 and no t is t^6
lcm of x^8 and x^3 is x^8
lcm of y^5 and y^6 is y^6
---
LCM = 12t^6*x^8*y^6
========================================================
3) Factor Completely: 27z^3+125
No common factor
-------------------------------
4) Factor: 9t+wt+9y+yw
---
t(9+w) + y(9+w)
---
= (9+w)(t+y)
-------------------------------

5) Factor: -8t+xt-24x+3x^2
---
= -(8t -xt + 24x - 3x^2)
---
= -(t(8-x) + 3x(8-x))
---
= -(8-x)(t+3x)
==================
Cheers,
Stan H.
==================

Answer by nyc_function(2741)

How about posting one question at a time? If you post more than one question, no one will respond.
Factor Completely: 27z^3+125
Use the sum of cubes formula: a^3 + b^3 = (a + b)(a^2 - 2b + b^2)
=================================================================
27z^3 + 125 can be written as (3z)^3 + 5^3.
(3z)^3 + 5^3 = (3z + 5)(3z^2 - 3z*5 + 5^2)
(3z)^3 + 5^3 = (3z + 5)(3z^2 -15z + 25)
Done!